OpenStudy (rainbow_rocks03):

Plz explain how to do this NO direct answers plz! will medal for best response.

OpenStudy (anonymous):

hi

OpenStudy (rainbow_rocks03):

i am putting the question down plz be patient, TY!

OpenStudy (rainbow_rocks03):

PLZ explain no direct answers!

OpenStudy (rainbow_rocks03):

@kimbllse777

OpenStudy (anonymous):

you need to change both fractions to a common denominator to do addition (and subtraction). What is the least common multiple of 6 and 8?

OpenStudy (rainbow_rocks03):

do you mean LCD or LCM cause its been saying LCD

OpenStudy (anonymous):

for this purpose, they're pretty much the same thing

OpenStudy (rainbow_rocks03):

oh ok, i thought so

OpenStudy (rainbow_rocks03):

So the LCD of 6 and 8 is 24?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

hi

OpenStudy (rainbow_rocks03):

stop saying hi plz.

OpenStudy (anonymous):

so now you have to multiply the fractions by 1 in other forms to change them. multiply 1/6 by 4/4 and mulitply 3/8 by 3/3.\[\frac{ 1\times4 }{ 6\times4 }+\frac{ 3\times3 }{ 8\times3 }\]

OpenStudy (anonymous):

that will give you 24 in both denominators

OpenStudy (rainbow_rocks03):

1*4=4 & 3*3=9

OpenStudy (anonymous):

yes, so you have 4/24 + 9/24 =

OpenStudy (rainbow_rocks03):

Now I add the numerators together?

OpenStudy (anonymous):

yes

OpenStudy (rainbow_rocks03):

so 4+9=13

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so your final answer is what?

OpenStudy (rainbow_rocks03):

13/24 ?

OpenStudy (anonymous):

yep :)

OpenStudy (rainbow_rocks03):

TY for the help @peachpi !!! :D

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

the reason this works is because once the two values are expressed with a common denominator, that fraction can be factored out. it leaves a simple addition of whole numbers which can be added and then multipled by the factor to yield a single value: \[\frac{ 1 }{ 6 }+\frac{ 3 }{ 8 }=\frac{ 1 }{ 6 }\cdot \frac{ 4 }{ 4 }+\frac{ 3 }{ 8 }\cdot \frac{ 3 }{ 3 }=\frac{ 4 }{ 24}+\frac{ 9 }{ 24 }=\frac{ 1 }{ 24 }\left( 4+9 \right)=\frac{ 1 }{ 24 }\left( 13 \right)=\frac{ 13 }{ 24 }\]

OpenStudy (rainbow_rocks03):

can you help with another cause im still kinda not used to it. :/

OpenStudy (rainbow_rocks03):

@peachpi ?